Grade 12Physical Chemistry

The decomposition of NH3 on platinum surface is zero order reaction . If rate constant (k) is 4*10^-3 , how long will it take to reduce the initial concentration of NH3 from 0.1 to 0.064 M

himanshu
7 Years agoGrade 12

1 Answer

Khimraj
7 Years ago
d[NH3]/dt = 4*10-3
t = 4*10-3/(0.1 – 0.064) sec = 1/9 sec
Hope it clears.